Part A
In late 2020, the Nicklaus Corporation was formed. The corporate charter authorizes the issuance of 4,000,000 shares of common stock carrying a $1 par value, and 1,000,000 shares of $5 par value, noncumulative, nonparticipating preferred stock. On January 2, 2021, 2,000,000 shares of the common stock are issued in exchange for cash at an average price of $12 per share. Also on January 2, all 1,000,000 shares of preferred stock are issued at $30 per share.

Required:
1. Prepare journal entries to record these transactions.
2. Prepare the shareholders' equity section of the Nicklaus balance sheet as of March 31, 2021. (Assume net income for the first quarter 2021 was $1,300,000.)

Part B
During 2021, the Nicklaus Corporation participated in three treasury stock transactions:

  1. On June 30, 2021, the corporation reacquires 150,000 shares for the treasury at a price of $14 per share.
  2. On July 31, 2021, 25,000 treasury shares are reissued at $17 per share.
  3. On September 30, 2021, 25,000 treasury shares are reissued at $12 per share.


Required:
1. Prepare journal entries to record these transactions.
2. Prepare the Nicklaus Corporation shareholders' equity section as it would appear in a balance sheet prepared at September 30, 2021. (Assume net income for the second and third quarter was $2,750,000.)

Part C
On October 1, 2021, Nicklaus Corporation receives permission to replace its $1 par value common stock (4,000,000 shares authorized, 2,000,000 shares issued, and 1,900,000 shares outstanding) with a new common stock issue having a $0.50 par value. Since the new par value is one-half the amount of the old, this represents a 2-for-1 stock split. That is, the shareholders will receive two shares of the $0.50 par stock in exchange for each share of the $1 par stock they own. The $1 par stock will be collected and destroyed by the issuing corporation.

On November 1, 2021, the Nicklaus Corporation declares a $0.09 per share cash dividend on common stock and a $0.26 per share cash dividend on preferred stock. Payment is scheduled for December 1, 2021, to shareholders of record on November 15, 2021.

On December 2, 2021, the Nicklaus Corporation declares a 3% stock dividend payable on December 28, 2021, to shareholders of record on December 14. At the date of declaration, the common stock was selling in the open market at $12 per share. The dividend will result in 114,000 (0.03 × 3,800,000) additional shares being issued to shareholders.

Required:
1. Prepare journal entries to record the declaration and payment of these stock and cash dividends.
2. Prepare the December 31, 2021, shareholders' equity section of the balance sheet for the Nicklaus Corporation. (Assume net income for the fourth quarter was $2,250,000.)
3. Prepare a statement of shareholders' equity for Nicklaus Corporation for 2021.Answer is not complete. Complete this question by entering your answers in the tabs below. Required 1 Part A Required 2 PartComplete this question by entering your answers in the tabs below. Required 1 Part A Required 2 Part A Required 1 Part B RequAnswer is not complete. Complete this question by entering your answers in the tabs below. Required 1 Part A Required 2 PartComplete this question by entering your answers in the tabs below. Required 1 Part A Required 2 Part A Required i Part B Requx Answer is not complete. Complete this question by entering your answers in the tabs below. Required 1 Part A Required 2 ParComplete this question by entering your answers in the tabs below. Required 1 Part A Required 2 Part A Required 1 Part B RequRequired 1 Part A Required 2 Part A Required 1 Part B Required 2 Part B Required 1 Part C Required 2 Part C Required 3 Part C
Answer #1

Part A

Shareholder's Equity
Common Stock $     20,00,000
Preferred Stock $     50,00,000
Paid in capital in excess of par $ 4,70,00,000
Retained Earnings $     13,00,000
Total Shareholder's Equity $ 5,53,00,000

Part B

Date Account Titles Debit Credit
Jun-30 Treasury Stock $     21,00,000
       Cash $   21,00,000
Jul-31 Cash $       4,25,000 =25000*17
       Treasury Stock     $     3,50,000 =25000*14
       Paid in capital from Treasury Stock $        75,000 =425000-350000
Sep-30 Cash $       3,00,000 =25000*12
Paid in capital from Treasury Stock $          50,000 =350000-300000
       Treasury Stock $     3,50,000 =25000*14
Shareholder's Equity
Common Stock $     20,00,000
Preferred Stock $     50,00,000
Paid in capital in excess of par $ 4,70,00,000
Paid in capital from Treasury Stock $          25,000
Retained Earnings $     13,00,000
$ 5,53,25,000
Less : Treasury Stock $     14,00,000
Total Shareholder's Equity $ 5,39,25,000

As per HOMEWORKLIB RULES we are suppose to answer 4 parts, i have answered upto first 4, so kindly post other parts separately